# Single-shot near-field reconstruction of metamaterial dispersion

### v2

@@ Abstract
We present a single-shot near-field technique to reconstruct the isofrequency surfaces of metamaterials in the microwave regime. In our approach, we excite resonant modes using a fixed source in a resonator composed of the material under test and map the in-plane field distribution with a movable probe. Applying a fast Fourier transform (FFT) to the measured field reveals the sample’s in-plane dispersion. By extending this analysis over multiple frequencies and comparing the results with Fabry–Pérot resonances, we retrieve the full three-dimensional dispersion relation. When we apply the method to a double non-connected wire metamaterial, it accurately captures the low-frequency hyperbolic isofrequency surface, providing both a precise experimental tool and conceptual insight into spatially dispersive metamaterials.
@@ I Introduction
The recent development of metamaterials has significantly broadened the scope for exploring novel wave propagation phenomena. This progress stems from the unprecedented design flexibility offered by meta-atoms—the fundamental building blocks of metamaterials. Metamaterials have enabled experimental demonstrations of negative refraction [shalaev2007optical], backward waves [silin2012history], and other unconventional electromagnetic effects [capolino2017theory]. Additionally, recent designs have exhibited exceptionally high effective permittivity along specific propagation directions [poddubny2013hyperbolic, enriquez_2024], surpassing the intrinsic limits of natural materials such as ceramics [kell1973high], oxides [taylor2019fundamental, peng2020origin], perovskites [peng2020new], and metal–dielectric composites in the low-frequency regime [du2016colossal]. In particular, hyperbolic metamaterials [poddubny2013hyperbolic]—a class of anisotropic media with permittivity tensor components of opposite sign—exhibit hyperbolic isofrequency contours. These materials support extreme electromagnetic responses across a wide frequency range, from microwaves to the optical domain.
The concept of dispersion provides a fundamental framework for describing wave phenomena in both continuous media and structured materials. Dispersion relations characterize how wave behavior depends on frequency and wave vector via the material’s effective parameters. Temporal dispersion refers to the frequency dependence of these parameters, while spatial dispersion accounts for nonlocal effects, in which the material response explicitly depends on the wave vector [landau1960course, born]. Analyzing a material’s dispersion relations offers valuable insight into the interaction of both propagating and evanescent waves with the structure.
Dispersion diagrams, isofrequency contours and surfaces have long been widely accepted as methods for describing and analyzing the electromagnetic properties of materials and wave propagation [joannopoulos2008molding, tretyakov2003analytical]. Analytical determination of these dispersion characteristics necessitates knowledge of the material parameter tensors [born]. For metamaterials, which are the focus of the present study, various homogenization methods can be used to determine the effective material parameters [alu2011first, silveirinha2007metamaterial]. However, such estimates often fail to be universal over a wide frequency range.
@@
Developing a general-purpose algorithm for the theoretical description of any metamaterial appears to be practically impossible due to the variety of metamaterials and complexities and limitations of the existing homogenization approaches. However, creating experimental samples of materials with specific properties is feasible and promising, even without a universal description. For instance, various metamaterial prototypes have been fabricated exhibiting hyperbolic isofrequency contours for applications in control of light propagation [poddubny2013hyperbolic].
Today, the dispersion of any arbitrary metamaterial can be easily calculated by solving the Maxwell equations using numerical methods [capolino2017theory]. In many cases, we do not even have to delve into the physical subtleties of the problem; it is enough to determine the unit cell and the materials constituting it. However, the experimental confirmation of the properties found in simulations or theory remains a very challenging task.
Researchers have studied the dispersion properties of planar structures in the optical regime using back focal plane microscopy, which enables reconstruction of isofrequency contours [permyakov_visualization_2018, pidgayko_direct_2019, jin_topologically_2019]. Isofrequency contours have also been obtained from numerical simulations of infinite cavity samples of different sizes, each resonating at the same frequency [yang2012experimental]. In the microwave range, several techniques have been used to reconstruct dispersion characteristics of structured materials, including coherent microwave transient spectroscopy [robertson_measurement_1992, tang_experimental_2023], near-field scanning combined with a fast Fourier transform (FFT) [lemoult_resonant_2010, yves_measuring_2018, tremain2018isotropic, hooper2014massively], and complex-field measurements, in which wave vectors are inferred from phase differences per unit length [belov2013single, itoh2005electromagnetic]. However, full three-dimensional dispersion properties—including isofrequency surfaces—have not yet been retrieved in a single measurement without rotating the sample.
Our study proposes an experimental approach that requires a single measurement of the electromagnetic near-field distribution of a multi-mode resonator. This method allows for comprehensive three-dimensional dispersion characterization across a wide frequency range by extracting isofrequency contours of several waveguide modes in the $k_{x}k_{y}$-plane. The $k_{z}$ component of each mode is identified via the Fabry-Pérot resonance condition along the $z$-direction. We showcase the effectiveness of our approach through an example wherein we reconstruct the three-dimensional isofrequency surfaces of the hyperbolic mode within the double non-connected wire metamaterial [simovski2004low].
@@ II Main concept
Figure 1(a) presents a schematic of the proposed three-dimensional dispersion extraction method. A source excites waves that propagate through the sample, reflect from its boundaries, and interfere within the enclosed volume. A movable probe scans the top xxyy plane, measuring the electromagnetic field distribution point by point. While the in-plane dispersion relation $\omega(k_{x},k_{y})$ can be obtained using a standard FFT technique [tremain2018isotropic], the out-of-plane wave-vector component $k_{z}$ is not known a priori. However, when the refractive index of the sample is much higher than that of air—or when the sample is bounded by metallic walls—the system supports resonant modes that satisfy the Fabry–Pérot condition $D_{i}k_{i}=n_{i}\pi$, where $i=x,y,z$, $D_{i}$ is the sample dimension along the $i$ direction, and $n_{i}$ is an integer representing the resonance order. This relation enables discrete sampling of the dispersion along $k_{z}$, since each spectral peak—where $k_{x}$ and $k_{y}$ are already determined—corresponds to a resonant mode with a specific $n_{z}$ value.
@@ III.1 Wire metamaterials
To illustrate the method in detail, we consider an artificial material composed of metallic wires. Among all metamaterials, the class of wire media is particularly notable for its pronounced spatial dispersion (nonlocality), which arises from the unrestricted flow of charges along the metallic wires. Although various types of wire-based metamaterials have been proposed, only specific subclasses have been rigorously described using an effective permittivity tensor. The simple wire medium [belov2003strong], as well as double and triple non-connected wire media [simovski2004low, sakhno2022quadraxial], belong to this category. In contrast, other wire metamaterials still lack a comprehensive analytical description.
The method of isofrequency surface reconstruction introduced here provides a powerful tool to bridge this gap. It serves a dual purpose: validating existing effective-medium models and corroborating numerical simulations of nonhomogenized metamaterials. In this work, we focus on the double non-connected wire medium, which exhibits a hyperbolic-like mode at low frequencies [simovski2004low].
@@ III.2 Double non-connected wire metamaterial
Figure 2: (a) Unit cell of the double non-connected wire metamaterial. (b) Isofrequency surface $\omega=0.2(2\pi c/a)$ (below the plasma frequency) plotted according to the effective permittivity tensor derived in [simovski2004low].
Based on the effective dielectric tensor derived in the low-frequency limit [simovski2004low], two modes emerge from zero frequency. The corresponding isofrequency surfaces, obtained by substituting the effective permittivity into the Fresnel equation, are shown in Fig. 2(b).
The first mode (black surface, Fig. 2(b)) is an ordinary mode, also known as the low-$k$ mode [simovski2004low], with TE polarization, where the electric field points along the $z$-axis [see Fig. 2(a)]. In the $k_{z}=0$ plane, its isofrequency surface forms a circular cross-section with a radius equal to the vacuum wave vector at the corresponding frequency. In cross-sections perpendicular to $k_{z}$, the in-plane wave vector projection remains close to the vacuum wave number, reflecting the negligible polarizability of the thin wires.
@@ III.4 Near-field scanning and FFT
In our experimental setup, the near-field distribution above the sample was measured. Owing to the polarization of the high-$k$ modes, the measured transmission parameter S21 is proportional to the $H_{z}$ field distribution at the corresponding frequency. Using a vector network analyzer (VNA), we recorded the transmission between two electrically small loop antennas over a broad frequency range [see Fig. 4(a,b)]. The source antenna was placed near the center of the resonator, inside the sample, to excite all supported TM modes through near-field coupling. The receiving antenna—also a small loop oriented to detect the magnetic field component normal to the sample surface—was positioned 3 mm above the sample.
Figure 4: Measurements of the $H_{z}$ field near the sample. (a) Transmission coefficient between the source and an arbitrarily positioned receiver; the numbers indicate the number of half-wavelengths across the wire array. (b) Frequency range covering all responses with the same $k_{y}=2\pi/(N_{y}a)$ order. (c) Examples of near-field distributions in real space for selected frequencies, along with the corresponding Fourier spectra.
@@
Although the probe-antenna method is inherently limited to the microwave regime, the underlying approach—reconstructing isofrequency surfaces from spatially resolved field measurements—is conceptually applicable to higher-frequency domains; in optics, for example, near-field scanning probes and Fourier/back-focal-plane or leakage-radiation imaging have been used to obtain $k$-space (isofrequency) maps and extract dispersion of guided and surface modes [Betzig1992, Regan2016, Vasista2019, Jang2024].
@@ IV Results and discussion
To investigate the properties of the waveguide modes, we performed eigenmode simulations using CST STUDIO SUITE [cst]. The numerical model implemented periodic boundary conditions in the wire plane ($x$-$y$), and absorbing boundary conditions (or Perfectly Matched Layers (PMLs)) were applied in the $z$ direction. The $k_{y}$ component was fixed to the experimentally determined value [see Fig. 5(a)], while the $k_{x}$ wave vector was swept across the first Brillouin zone, ranging from 0 to $\pi/a\,\mathrm{m}^{-1}$. The resulting dispersion diagram, shown in Fig. 5(b), reveals three distinct branches: the fundamental mode, $\mathrm{WG}_{0}$ [Fig. 5(c)], and a pair of high-order modes [Figs. 5(d) and 5(e)].
@@
To extract the dispersion branches from a specific $k_{y}$-slice of the spectrum map [Fig. 6(a)], we manually select the three sets of bright points and then apply a least-squares interpolation to generate smooth dispersion curves, as shown in Fig. 6(b). Interestingly, researchers in geophysics face a similar challenge when separating dispersion branches and have addressed it using deep learning methods [dong2021dispernet, dai2021deep].
@@
We reconstruct the isofrequency surfaces from the extracted two-dimensional dispersion $f(k_{x},k_{y})$ [Fig. 6(c)] for each mode corresponding to $k_{z}=n_{z}\pi/(aN_{z})$. The surfaces for selected frequencies below the plasma frequency of the wire-medium sample ($f=4,8,12~\mathrm{GHz}$) are shown in Fig. 7. Figures 7(a–c) present the analytically obtained isofrequency surfaces [simovski2004low], while Figs. 7(d–f) display the corresponding experimentally measured contours (red lines) together with the numerical results obtained from the CST eigenfrequency solver [cst] (gray surfaces). The experimentally reconstructed contours show excellent agreement with both the analytical and numerical results, reproducing the hyperbolic behavior in the $k_{x}k_{y}$ plane expected for the TM mode.
@@ V Conclusion
We have introduced a method to experimentally reconstruct isofrequency surfaces in three-dimensional metamaterials. The method relies on a single-shot near-field scan of the resonant modes of a resonator composed of the target metamaterial. By applying fast Fourier transforms over a wide frequency range and comparing the observed resonances with a Fabry–Pérot model, we obtain the isofrequency surfaces. We have demonstrated the method’s effectiveness using a resonator based on a double non-connected wire metamaterial, confirming the hyperbolic isofrequency surface of one of its two low-frequency modes. The approach is straightforward to implement and enables rapid, comprehensive dispersion characterization. Although it requires samples with a refractive index significantly higher than that of air, it can be extended to more general configurations where electric walls bound the structure; in such cases, field scans can be performed through predrilled holes in the cavity, as shown in Ref. [enriquez_2025].

### v3

@@ Abstract
We present a single-shot near-field technique, where the near-field scan is performed on a single sample without repeating measurements or averaging over multiple samples, to reconstruct the isofrequency surfaces of metamaterials in the microwave regime. In our approach, we excite resonant modes using a fixed source in a resonator composed of the material under test and map the in-plane field distribution with a movable probe. Applying a fast Fourier transform (FFT) to the measured field reveals the sample’s in-plane dispersion. By extending this analysis over multiple frequencies and comparing the results with Fabry–Pérot resonances, we retrieve the full three-dimensional dispersion relation. When we apply the method to a double non-connected wire metamaterial, it accurately captures the low-frequency hyperbolic isofrequency surface, providing both a precise experimental tool and conceptual insight into spatially dispersive metamaterials.
@@ I Introduction
The recent development of metamaterials has significantly broadened the scope for exploring novel wave propagation phenomena. This progress stems from the unprecedented design flexibility offered by meta-atoms—the fundamental building blocks of metamaterials. Metamaterials have enabled experimental demonstrations of negative refraction [30], backward waves [31], and other unconventional electromagnetic effects [6]. Additionally, recent designs have exhibited exceptionally high effective permittivity along specific propagation directions [26, 12], surpassing the intrinsic limits of natural materials such as ceramics [18], oxides [36, 23], perovskites [22], and metal–dielectric composites in the low-frequency regime [10]. In particular, hyperbolic metamaterials [26]—a class of anisotropic media with permittivity tensor components of opposite sign—exhibit hyperbolic isofrequency contours. These materials support extreme electromagnetic responses across a wide frequency range, from microwaves to the optical domain.
The concept of dispersion provides a fundamental framework for describing wave phenomena in both continuous media and structured materials. Dispersion relations characterize how wave behavior depends on frequency and wave vector via the material’s effective parameters. Temporal dispersion refers to the frequency dependence of these parameters, while spatial dispersion accounts for nonlocal effects, in which the material response explicitly depends on the wave vector [20, 5]. Analyzing a material’s dispersion relations offers valuable insight into the interaction of both propagating and evanescent waves with the structure.
Dispersion diagrams, isofrequency contours and surfaces have long been widely accepted as methods for describing and analyzing the electromagnetic properties of materials and wave propagation [17, 38]. Analytical determination of these dispersion characteristics necessitates knowledge of the material parameter tensors [5]. For metamaterials, which are the focus of the present study, various homogenization methods can be used to determine the effective material parameters [1, 33]. However, such estimates often fail to be universal over a wide frequency range.
@@
Developing a general-purpose algorithm for the theoretical description of arbitrary metamaterials is practically impossible due to their variety and the limitations of existing homogenization approaches. On the other hand, the dispersion of a specific metamaterial can often be readily calculated by solving Maxwell’s equations numerically [6]. In many cases, it is sufficient to know the unit cell geometry and the constituent materials, without delving into all the physical subtleties. However, while experimental prototypes have been fabricated [26], experimentally verifying the dispersion properties predicted by theory or simulations remains a challenging task.
Researchers have studied the dispersion properties of planar structures in the optical regime using back focal plane microscopy, which enables reconstruction of isofrequency contours [24, 25, 16]. Isofrequency contours have also been obtained from numerical simulations of infinite cavity samples of different sizes, each resonating at the same frequency [40]. In the microwave range, several techniques have been used to reconstruct dispersion characteristics of structured materials, including coherent microwave transient spectroscopy [28, 35], near-field scanning combined with a fast Fourier transform (FFT) [21, 41, 37, 13], and complex-field measurements, in which wave vectors are inferred from phase differences per unit length [3, 14]. However, full three-dimensional dispersion properties—including isofrequency surfaces—have not yet been retrieved in a single measurement without rotating the sample.
Our study proposes an experimental approach that requires a single measurement of the electromagnetic near-field distribution of a multi-mode resonator. This method allows for comprehensive three-dimensional dispersion characterization across a wide frequency range by extracting isofrequency contours of several waveguide modes in the $k_{x}k_{y}$-plane. The $k_{z}$ component of each mode is identified via the Fabry-Pérot resonance condition along the $z$-direction. We showcase the effectiveness of our approach through an example wherein we reconstruct the three-dimensional isofrequency surfaces of the hyperbolic mode within the double non-connected wire metamaterial [34].
@@ II Main concept
Figure 1(a) presents a schematic of the proposed three-dimensional dispersion extraction method. A source excites waves that propagate through the sample, reflect from its boundaries, and interfere within the enclosed volume. A movable probe scans the top xxyy plane, measuring the electromagnetic field distribution point by point. While the in-plane dispersion relation $\omega(k_{x},k_{y})$ can be obtained using a standard FFT technique [37], the out-of-plane wave-vector component $k_{z}$ is not known a priori. However, when the refractive index of the sample is much higher than that of air—or when the sample is bounded by metallic walls—the system supports resonant modes that satisfy the Fabry–Pérot condition $D_{i}k_{i}=n_{i}\pi$, where $i=x,y,z$, $D_{i}$ is the sample dimension along the $i$ direction, and $n_{i}$ is an integer representing the resonance order. This relation enables discrete sampling of the dispersion along $k_{z}$, since each spectral peak—where $k_{x}$ and $k_{y}$ are already determined—corresponds to a resonant mode with a specific $n_{z}$ value.
@@ III.1 Wire metamaterials
To illustrate the method in detail, we consider an artificial material composed of metallic wires. Among all metamaterials, the class of wire media is particularly notable for its pronounced spatial dispersion (nonlocality), which arises from the unrestricted flow of charges along the metallic wires. Although various types of wire-based metamaterials have been proposed, only specific subclasses have been rigorously described using an effective permittivity tensor. The simple wire medium [2], as well as double and triple non-connected wire media [34, 29], belong to this category. In contrast, other wire metamaterials still lack a comprehensive analytical description.
The method of isofrequency surface reconstruction introduced here provides a powerful tool to bridge this gap. It serves a dual purpose: validating existing effective-medium models and corroborating numerical simulations of nonhomogenized metamaterials. In this work, we focus on the double non-connected wire medium, which exhibits a hyperbolic-like mode at low frequencies [34].
@@ III.2 Double non-connected wire metamaterial
Figure 2: (a) Unit cell of the double non-connected wire metamaterial. (b) Isofrequency surface $\omega=0.2(2\pi c/a)$ (below the plasma frequency) plotted according to the effective medium model of the metamaterial
[34].
Solving the system of equations (40) from [34] with the effective permittivity tensor of the metamaterial substituted (see Eqs. (42–43) of [34]) leads to Eq. (36) of the same reference. At a fixed frequency, the determinant of the system (40) vanishes (which is equivalent to satisfying Eq. (36)) for the eigen wave vectors of the metamaterial. Consequently, the isofrequency surfaces can be obtained and plotted in reciprocal space. In Fig. 2(b), we present analytical isofrequency surfaces for the homogenized [2, 32, 33] double non-connected wire metamaterial below its plasma frequency.
The first mode (black surface, Fig. 2(b)) is an ordinary mode, also known as the low-$k$ mode [34], with TE polarization, where the electric field points along the $z$-axis [see Fig. 2(a)]. In the $k_{z}=0$ plane, its isofrequency surface forms a circular cross-section with a radius equal to the vacuum wave vector at the corresponding frequency. In cross-sections perpendicular to $k_{z}$, the in-plane wave vector projection remains close to the vacuum wave number, reflecting the negligible polarizability of the thin wires.
@@ III.4 Near-field scanning and FFT
Figure 4: Measurements of the $H_{z}$ field near the sample. (a) Transmission coefficient between the source located and an arbitrarily positioned receiver; the numbers indicate the number of half-wavelengths across the wire array. (b) Frequency range covering all responses with the same $k_{y}=2\pi/(N_{y}a)$ order. (c) Examples of near-field distributions in real space for selected frequencies (red dots in panel b), along with the corresponding Fourier spectra.
@@
In our experimental setup, the near-field distribution above the sample was measured. Owing to the polarization of the high-$k$ modes, the measured transmission parameter S21 is proportional to the $H_{z}$ field distribution at the corresponding frequency. Using a vector network analyzer (VNA), we recorded the transmission between two electrically small loop antennas over a broad frequency range [see Fig. 4(a,b)]. The source antenna was placed near the center of the resonator, inside the sample, to excite all supported TM modes through near-field coupling. The receiving antenna—also a small loop oriented to detect the magnetic field component normal to the sample surface—was positioned 3 mm above the sample. For the spectra presented in Fig. 4(a,b), the receiver was placed at a position offset by fifteen unit cells in both the vertical and horizontal directions from the center of the sample, as indicated in the top-view inset of Fig. 4(a).
@@
Although the probe-antenna method is inherently limited to the microwave regime, the underlying approach—reconstructing isofrequency surfaces from spatially resolved field measurements—is conceptually applicable to higher-frequency domains; in optics, for example, near-field scanning probes and Fourier/back-focal-plane or leakage-radiation imaging have been used to obtain $k$-space (isofrequency) maps and extract dispersion of guided and surface modes [4, 27, 39, 15].
@@ IV Results and discussion
To investigate the properties of the waveguide modes, we performed eigenmode simulations using CST STUDIO SUITE [8]. The numerical model implemented periodic boundary conditions in the wire plane ($x$-$y$), and absorbing boundary conditions (or Perfectly Matched Layers (PMLs)) were applied in the $z$ direction. The $k_{y}$ component was fixed to the experimentally determined value [see Fig. 5(a)], while the $k_{x}$ wave vector was swept across the first Brillouin zone, ranging from 0 to $\pi/a\,\mathrm{m}^{-1}$. The resulting dispersion diagram, shown in Fig. 5(b), reveals three distinct branches: the fundamental mode, $\mathrm{WG}_{0}$ [Fig. 5(c)], and a pair of high-order modes [Figs. 5(d) and 5(e)].
@@
To extract the dispersion branches from a specific $k_{y}$-slice of the spectrum map [Fig. 6(a)], we manually select the three sets of bright points and then apply a least-squares interpolation to generate smooth dispersion curves, as shown in Fig. 6(b). Interestingly, researchers in geophysics face a similar challenge when separating dispersion branches and have addressed it using deep learning methods [9, 7].
@@
We reconstruct the isofrequency surfaces from the extracted two-dimensional dispersion $f(k_{x},k_{y})$ [Fig. 6(c)] for each mode corresponding to $k_{z}=n_{z}\pi/(aN_{z})$. The surfaces for selected frequencies below the plasma frequency of the wire-medium sample ($f=4,8,12~\mathrm{GHz}$) are shown in Fig. 7. Figures 7(a–c) present the analytically obtained isofrequency surfaces [34], while Figs. 7(d–f) display the corresponding experimentally measured contours (red lines) together with the numerical results obtained from the CST eigenfrequency solver [8] (gray surfaces). The experimentally reconstructed contours show excellent agreement with both the analytical and numerical results, reproducing the hyperbolic behavior in the $k_{x}k_{y}$ plane expected for the TM mode.
@@ V Conclusion
We have introduced a method to experimentally reconstruct isofrequency surfaces in three-dimensional metamaterials. The method relies on a single-shot near-field scan of the resonant modes of a resonator composed of the target metamaterial. By applying fast Fourier transforms over a wide frequency range and comparing the observed resonances with a Fabry–Pérot model, we obtain the isofrequency surfaces. We have demonstrated the method’s effectiveness using a resonator based on a double non-connected wire metamaterial, confirming the hyperbolic isofrequency surface of one of its two low-frequency modes. The approach is straightforward to implement and enables rapid, comprehensive dispersion characterization. Although it requires samples with a refractive index significantly higher than that of air, it can be extended to more general configurations where electric walls bound the structure; in such cases, field scans can be performed through predrilled holes in the cavity, as shown in Ref. [11].
Acknowledgements.
The work is funded by Russian Science Foundation grant No. 25-12-00261 (https://rscf.ru/project/25-12-00261/).
Data availability
The data that support the findings of this article are openly available [19].